# Hierarchical Uncertainty-Aware Graph Neural Network

### v1

@@ 4.2. Community Pooling
After one or more local message‐passing layers (we can stack several if needed), HU‑GNN elevates the node‑level representations to the group level. As shown in Figure 1, we introduce a node for each community $C_{m}$ (depicted as squares). First, we determine each node’s community membership via a trainable matrix as,
(8) $\underbrace{p_{i\to C_{m}}}_{\text{score for }C_{m}}=\frac{\exp\bigl{(}w_{m}^{\top}h_{i}\bigr{)}}{\sum_{m^{\prime}=1}^{m}\exp\bigl{(}w_{m^{\prime}}^{\top}h_{i}\bigr{)}},$
where the assignment scores for node $i$ are produced by a row‑wise softmax over the learnable weight matrix $W_{M}=[\mathbf{w}_{1},\dots,\mathbf{w}_{m}]^{\top}$. We set $M$ equal to the number of classes in the dataset. Then the community assignment process is given by:
(9) $i_{c}=\operatorname*{arg\,max}\limits_{m=1,\dots,M}\;p_{i\to C_{m}},$
@@
The discrete community assignment $i_{c}$ is used for the pooling step. Each community’s representation is computed by aggregating the embeddings of its member nodes. Accordingly, the community aggregation can be expressed as:
(10) ${h}_{C_{m}}=\frac{1}{|C_{m}|}\sum_{i=1}^{N}\mathbf{1}(i_{c}=m)\;W_{C}\,h_{i},$
@@
which is a simple mean-pooling of node features $h_{i}$ that belong to the same community $C_{m}$. The $W_{C}$ is the learnable weight matrix for this process. At the same time, we compute the community uncertainty $u_{C_{m}}$ from the dispersion of its member embeddings. Specifically, the variance of the member features around the community mean reflects how cohesive that community is:
(11) $u_{C_{m}}=f_{u}\Biggl{(}\frac{1}{\lvert C_{m}\rvert}\sum_{i\in C_{m}}\bigl{\|}h_{i}-h_{C_{m}}\bigr{\|}^{2}\Biggr{)},$
@@ 4.3. Global Integration
We introduce a global node $G$ (distinct from the graph $\mathcal{G}$) that is connected to all community nodes. We regard $h_{G}$ as a global node representation. Here, we define global integration as follows:
(12) ${h}_{G}=\frac{1}{M}\sum_{m=1}^{M}{W}_{G}{h}_{C_{m}}$
@@
(13) $u_{G}=f_{u}\left(\frac{1}{M}\sum_{m=1}^{M}\|{h}_{G}-{h}_{C_{m}}\|^{2}\right)$
@@ 4.4. Inference
(14) ${h}_{i}^{\text{final}}=\lambda_{i}{h}_{i}\;+\;\lambda_{C}{h}_{C_{m}}\;+\;\lambda_{G}{h}_{G},$
For each $v,w\in\{i,\,C_{m},\,G\}$, we compute the contribution weight $\lambda_{v}$ (where $\lambda_{i}$, $\lambda_{C}$, and $\lambda_{G}$ correspond to the local, community, and global terms, respectively) as,
(15) $\lambda_{v}\;=\;\frac{1}{2}\left(\frac{\exp(a_{v})}{\sum_{w}\exp(a_{w})}+\frac{\exp\bigl{(}-u_{v}\bigr{)}}{\sum_{w}\exp\bigl{(}-u_{w}\bigr{)}}\right)$
where $\alpha_{v}=\vec{a}^{T}\left(h_{i}\;\|\;h_{v}\right)$. The $\lambda$ balances how much a node trusts its estimation compared to the community’s uncertainty and feature similarity. In heterophilic cases, often $u_{i}$ is high since its neighbors were confusing. However, $u_{C_{m}}$ might be lower if the community contains some far-away same-class nodes that agreed. Thus, this update could significantly reduce uncertainty for such nodes, effectively stabilizing the prediction in heterophilic settings. Conversely, if the community is as uncertain as the node, nothing changes. Given the final representation of each node ${h}_{i}^{\text{final}}$, we pass them to a classifier (e.g., single-layer network) to predict the class probabilities for node $i$ as below:
@@
2:Label prediction of node $i$ ($\hat{y}_{i}$)
@@
5:for $\ell\leftarrow 1$ to $L$ do
6:     $\tilde{h}^{(\ell)}_{i}={W}_{O}^{(\ell)}{h}_{i}^{(\ell-1)}$
7:     $\alpha^{(\ell)}_{ij}\leftarrow\vec{a}^{T}\left(\tilde{h}^{(\ell)}_{i}\;\|\;\tilde{h}^{(\ell)}_{i}\right)$
8:     $u_{i}^{(\ell)}\leftarrow f_{u}\Bigl{(}\tfrac{1}{|N(i)|}\sum_{j\in N(i)}\|\,\tilde{h}^{(\ell)}_{i}-\tilde{h}^{(\ell)}_{j}\,\|^{2}\Bigr{)}$
9:     $m_{ij}^{(\ell)}\leftarrow\frac{1}{2}\left(\frac{\exp(a_{ij}^{(\ell-1)})}{\sum_{k\in N(i)}\exp(a_{ik}^{(\ell-1)})}+\frac{\exp(-u_{j}^{(\ell-1)})}{\sum_{k\in N(i)}\exp(-u_{k}^{(\ell-1)})}\right)$
10:     $h_{i}^{(\ell)}\leftarrow\sigma\Bigl{(}\tilde{h}^{(\ell)}_{i}+\sum_{j\in N(i)}m_{ij}^{(\ell)}\,\tilde{h}^{(\ell)}_{j}\Bigr{)}$
11$\triangleright$ Community Pooling (Sec. 4.2)
12:Assume $i\in C_{1}$
13:$h_{C_{1}}\leftarrow\frac{1}{|C_{1}|}\sum_{i=1}^{N}\mathbf{1}(i_{c}=1)\;W_{C}\,h_{i}$
14:$u_{C_{1}}\leftarrow f_{u}\Bigl{(}\tfrac{1}{|C_{1}|}\sum_{i\in C_{1}}\|\,h_{i}-h_{C_{1}}\|^{2}\Bigr{)}$
15$\triangleright$ Global Integration (Sec. 4.3)
16:$h_{G}\leftarrow\frac{1}{M}\sum_{m=1}^{M}{W}_{G}{h}_{C_{1}}$
17:$u_{G}\leftarrow f_{u}\left(\frac{1}{M}\sum_{m=1}^{M}\|{h}_{G}-{h}_{C_{1}}\|^{2}\right)$
18$\triangleright$ Attention and Inference (Sec. 4.4)
19:Compute $\lambda_{i}\;,\;\lambda_{C}\;,\;\lambda_{G}$
20:$h_{i}^{\mathrm{final}}\leftarrow\lambda_{i}\,h_{i}\;+\;\lambda_{C}\,h_{C_{1}}\;+\;\lambda_{G}\,h_{G}$
21:$\hat{y}_{i}\leftarrow\sigma\Bigl{(}W_{F}\,h_{i}^{\mathrm{final}}\Bigr{)}$
@@ 4.5. Optimization and Training Details
Initialization of Uncertainty.
For each node, we set the initial uncertainty $u_{i}^{(0)}$ according to the intrinsic reliability of its features. Concretely, we pre–compute a scalar confidence score with a node‑independent classifier $f_{\text{mlp}}(\cdot)$, a two–layer MLP with softmax trained on the raw attributes ${x}_{i}$. The initialization is given by:
(17) $u_{i}^{(0)}\;=\;1-\operatorname*{arg\,max}\;f_{\text{mlp}}({h}_{i}^{(0)})$
which means highly discriminative features ($u^{(0)}_{i}\approx 0$) begin with low uncertainty.
@@
    (19) $\mathcal{L}_{\text{sharp}}=\frac{1}{\lvert\mathcal{V}_{L}\rvert}\sum_{i\in\mathcal{V}_{\text{L}}}\mathbf{1}\bigl{(}\hat{y}_{i}=y_{i}\bigr{)}\,u_{i}$
@@
The hyper‑parameters $\beta_{1}$ and $\beta_{2}$ trade off supervision against calibration, which can be tuned using a validation set. Alternatively, when a fully probabilistic treatment is desired, the last two terms may be replaced with a Bayesian regularizer such as an evidence lower bound (ELBO) or a PAC‑Bayesian bound.
@@ 4.7. Computational Cost
Assume a graph $G=(\mathcal{V},\mathcal{E})$ with $|\mathcal{V}|=n$, $|\mathcal{E}|=m$, feature dimension $d$, $L$ local message‑passing layers, and $M$ communities. Each local layer $\ell$ first projects all $n$ node features in $O(n\,d^{2})$, then computes attention and uncertainty weights over the $m$ edges in $O(m\,d)$. Then, it aggregates neighbor messages in $O(m\,d)$, and updates node uncertainties in $O(m\,d)$ for a per‑layer cost of $O(n\,d^{2}+m\,d)$. Thus, the local cost follows $O\bigl{(}L\,(n\,d^{2}+m\,d)\bigr{)}$. Community pooling requires $O(n\,M\,d)$ to compute assignment scores for $n$ nodes across $M$ communities, $O(n\,M)$ for the $\arg\max$ assignments, and $O(n\,d)$ for mean‑pooling and variance computation $O(n\,M\,d)$. Finally, global integration and inference incur $O(M\,d^{2})$ to form the global embedding, $O(M\,d)$ to compute the global uncertainty, and $O(n\,d)$ to combine the final representations for all nodes, totaling $O(M\,d^{2}+n\,d)$. Altogether, the overall time complexity is $O\left(L\,(n\,d^{2}+m\,d)+n\,M\,d+M\,d^{2}+n\,d\right)$.
@@ 5. Theoretical Analysis
We now present theoretical results that characterize the performance of HU-GNN. We focus on three aspects: generalization ability, convergence of the uncertainty propagation mechanism, and robustness in heterophilic graphs. Proof sketches or intuitions are provided for each, with full proofs expected to be included in an appendix in a complete report.
@@ 5.1. PAC-Bayes Generalization Bounds
One of the advantages of modeling uncertainty is the potential to improve generalization by avoiding overfitting to noisy signals. We derive a PAC-Bayesian generalization bound for HU-GNN, showing that it achieves a tighter dependency on graph complexity measures (such as node degrees) compared to a standard GNN. Our analysis is inspired by the PAC-Bayes bounds for GNNs (Liao et al., 2020), which showed that for a GCN (Kipf and Welling, 2016), the maximum node degree $\Delta_{\max}$ and weight norms control the generalization gap. Intuitively, high-degree nodes are problematic because they can aggregate many noisy signals, increasing variance.
@@
Theorem 1 (PAC-Bayes Generalization Bound for HU-GNN) Consider the HU-GNN model with $L$ layers (including hierarchical ones) trained on a graph $G$ for node classification. Let $\mathcal{D}$ be the data distribution and $\mathcal{D}_{L}$ the training set (labels observed). Let $H[\cdot]$ denote the (empirical) entropy of predictions. Given that the loss function is bounded, the following bound holds with high probability for all posterior distributions $Q$ over the HU-GNN’s weights and a suitably chosen prior $P$:
@@
(21) $\begin{gathered}\mathbb{E}_{Q}[\text{Err}(\mathcal{D})]\;\leq\;\mathbb{E}_{Q}[\text{Err}(\mathcal{D}_{L})]\;+\;\\ \mathcal{O}\Big{(}\frac{1}{|\mathcal{D}_{L}|}\Big{[}\sum_{i\in\mathcal{D}_{L}}\min\{\tilde{\Delta}_{i},\Delta_{\max}\}\,+\,H[u^{(L)}]\Big{]}\Big{)}\;+\;\delta,\end{gathered}$
@@
, where $\text{Err}(\mathcal{D}_{L})$ is training error. Notation $\tilde{\Delta}_{i}$ is an effective degree of node $i$ after uncertainty-based reweighting, $H[u^{(L)}]$ is a measure of uncertainty entropy across the graph, and symbol $\delta$ is complexity term of order $\tilde{\mathcal{O}}\big{(}\frac{KL(Q\|P)+\log(1/\delta)}{|\mathcal{D}_{L}|}\big{)}$ for confidence $1-\delta$. In particular, $\min\{\tilde{\Delta}_{i},\Delta_{\max}\}$ indicates that HU-GNN effectively caps the influence of high-degree nodes by reducing weights of edges from uncertain neighbors.
@@
The precise form of the bound is technical, but the key insight is that HU-GNN’s uncertainty mechanism leads to an effective degree $\tilde{\Delta}_{i}$, which is often much smaller than the raw degree $\Delta_{i}$. For example, if node $i$ has 100 neighbors but 90 of them are deemed highly uncertain, the effective degree in the model’s hypothesis space is closer to 10. This means the model is less likely to overfit based on those 90 noisy neighbors. As a result, our PAC-Bayesian bound does not blow up with the actual maximum degree of the graph, but rather with a lower quantity reflecting the filtered graph connectivity. This yields a tighter generalization bound compared to a standard GNN on the original graph (Liao et al., 2020). In addition, the newly added term $H[u^{(L)}]$ penalizes high uncertainty entropy. This makes the bound looser if the model remains very unsure (high entropy in uncertainties) across the graph. Minimizing this term essentially encourages the model to explain away uncertainty when possible, which aligns with training objectives that reduce uncertainty as confidence improves. The detailed proof uses a PAC-Bayes bound on a Gibbs classifier that samples a random instantiation of the GNN weights, and leverages the convexity of the loss and the uncertainty gating to bound the change in loss when an edge weight is reduced. We also extend the perturbation analysis of message-passing networks to account for the random masking of edges by uncertainty, showing this acts like an $\ell_{0},\ell_{1}$ regularization on the adjacency. In conclusion, we prove that HU-GNN is theoretically justified to generalize better, especially on graphs with noisy connections.

### v2

@@ 4.2. Community Assignment and Pooling
After local message‑passing layers (we can stack more if needed), HU‑GNN elevates the node‑level representations to the group level. As shown in Figure 1, we softly assign each node to the community set $\{C_{1},\dots,C_{m}\}$ (depicted as squares).
Trainable assignment scores.
We first compute a score vector for node $i$ with a learnable weight matrix $W_{M}=[w_{1},\dots,w_{m}]^{\top}$:
(9) $p_{i\to C_{m}}\;=\;\frac{\exp\bigl{(}w_{m}^{\top}h_{i}\bigr{)}}{\sum_{j=1}^{m}\exp\bigl{(}w_{j}^{\top}h_{i}\bigr{)}},\qquad m=1,\dots,M.$
Differentiable community selection.
To make this process differentiable, we sample a one‑hot assignment $z_{i}\in\{0,1\}^{m}$ using the Gumbel‑Softmax estimator (Jang et al., 2016):
(10) $\tilde{a}_{i,m}=\operatorname{softmax}\Bigl{(}\bigl{(}\log p_{i\to C_{m}}+g_{m}\bigr{)}/\tau\Bigr{)},\quad g_{m}\sim\mathrm{Gumbel}(0,1),$
Here, $\tau>0$ is a temperature parameter that is annealed from 1.0 to 0.1 during training. During backpropagation, we employ the straight-through estimator. We define the community indicator as,
(11) $z_{i,m}=\begin{cases}1,&\text{if }m=\arg\max_{k}\tilde{a}_{i,k},\\ 0,&\text{otherwise,}\end{cases}$
@@
which is then used for the pooling step. We assume that $C_{m}=\{\,j\mid z_{j,m}=1\}$. Each community’s representation is obtained by aggregating the embeddings of its member nodes as below:
(12) ${h}_{C_{m}}=\sum_{j\in C_{m}}W_{C}\,h_{j},$
@@
which is a simple mean-pooling of node features $h_{i}$ that belong to the same community $C_{m}\in\{C_{1},\dots,C_{M}\}$ using the learnable matrix $W_{C}$. We also compute the uncertainty of the community $u_{C_{m}}$ from the spread of its member embeddings. Here, we use the variance of the member features in the same community to reflect cohesiveness:
(13) $u_{C_{m}}=f_{u}\Biggl{(}\frac{1}{|C_{m}|}\sum_{j\in C_{m}}\bigl{\|}h_{j}-h_{C_{m}}\bigr{\|}^{2}\Biggr{)},$
@@ 4.3. Global Integration
We introduce a global node $G$ (distinct from the graph $\mathcal{G}$) derived from all community nodes. We regard $h_{G}$ as a global node representation. Here, we define global integration as follows:
(14) ${h}_{G}=\frac{1}{m}\sum_{j=1}^{m}{W}_{G}{h}_{C_{j}}$
@@
(15) $u_{G}=f_{u}\left(\frac{1}{m}\sum_{j=1}^{m}\|{h}_{G}-{h}_{C_{j}}\|^{2}\right)$
@@ 4.4. Inference Stage
(16) ${h}_{i}^{\text{final}}=\lambda_{i}{h}_{i}\;+\;\lambda_{C_{m}}{h}_{C_{m}}\;+\;\lambda_{G}{h}_{G},$
For each $v,w\in\{i,\,C_{m},\,G\}$, we compute the contribution weight $\lambda_{v}$ (where $\lambda_{i}$, $\lambda_{C_{m}}$, and $\lambda_{G}$ correspond to the local, community, and global terms, respectively) as,
(17) $\lambda_{v}\;=\;\frac{\exp(a_{v})\exp\bigl{(}-u_{v}\bigr{)}}{\sum_{w}\exp(a_{w})\exp\bigl{(}-u_{w}\bigr{)}}$
where $\alpha_{v}=\vec{a}^{T}[h_{i}\;\|\;h_{v}]$ and $h_{v}\in\{h_{i},\,h_{C_{m}},\,h_{G}\}$. The $\lambda_{v}$ balances how much a node trusts its estimation compared to the community’s uncertainty and feature similarity. In heterophilic cases, $u_{i}$ is often high since its neighbors are confusing. However, $u_{C_{m}}$ might be lower if the community contains some far-away same-class nodes that agreed. Thus, this update could significantly reduce uncertainty for such nodes, effectively stabilizing the prediction in heterophilic settings. Conversely, nothing changes if the community is as uncertain as the node. Given the final representation of each node ${h}_{i}^{\text{final}}$, we pass them to a classifier (e.g., single-layer network) to predict the class probabilities for node $i$ as below:
@@
2:Predicted label of node $i$ ($\hat{y}_{i}$)
@@
5:for $\ell=1$ to $L$ do
6:     $\tilde{h}_{i}^{(\ell)}\leftarrow W_{O}^{(\ell)}\,h_{i}^{(\ell-1)}$
7:     $u_{i}^{(\ell)}\leftarrow f_{u}\!\bigl{(}\tfrac{1}{|N(i)|}\sum_{j\in N(i)}\|\tilde{h}_{i}^{(\ell)}-\tilde{h}_{j}^{(\ell)}\|^{2}\bigr{)}$
8:     $m_{ij}^{(\ell)}\leftarrow\dfrac{\exp\bigl{(}a^{\top}[\tilde{h}_{i}^{(\ell)}\|\tilde{h}_{j}^{(\ell)}]\bigr{)}\,\exp\bigl{(}-u_{j}^{(\ell)}\bigr{)}}{\sum_{k\in N(i)}\exp\bigl{(}a^{\top}[\tilde{h}_{i}^{(\ell)}\|\tilde{h}_{k}^{(\ell)}]\bigr{)}\,\exp\bigl{(}-u_{k}^{(\ell)}\bigr{)}}$
9:     $l_{i}^{(\ell)}\leftarrow\sum_{j\in N(i)}m_{ij}^{(\ell)}\,\tilde{h}_{j}^{(\ell)}$,    $d_{i}^{(\ell)}\leftarrow\tilde{h}_{i}^{(\ell)}-l_{i}^{(\ell)}$
10:     $h_{i}^{(\ell)}\leftarrow\sigma\Bigl{(}\tilde{h}_{i}^{(\ell)}+p^{(\ell)}_{i}\,l_{i}^{(\ell)}\,+(1-p^{(\ell)}_{i})\,d_{i}^{(\ell)}\Bigr{)}$
11:           $\diamond\;p^{(\ell)}_{i}=\cos\bigl{(}\tilde{h}_{i}^{(\ell)},\,l_{i}^{(\ell)}\bigr{)}$
12$\triangleright$ Community Assignment and Pooling (Sec. 4.2)
13:Assume $i\in C_{1}$
14:$h_{C_{1}}\leftarrow\frac{1}{|C_{1}|}\sum_{j\in C_{1}}W_{C}\,h_{j}$
15:$u_{C_{1}}\leftarrow f_{u}\Bigl{(}\tfrac{1}{|C_{1}|}\sum_{j\in C_{1}}\|\,h_{j}-h_{C_{1}}\|^{2}\Bigr{)}$
16$\triangleright$ Global Integration (Sec. 4.3)
17:$h_{G}\leftarrow\frac{1}{m}\sum_{j=1}^{m}{W}_{G}{h}_{C_{j}}$
18:$u_{G}\leftarrow f_{u}\left(\frac{1}{m}\sum_{j=1}^{m}\|{h}_{G}-{h}_{C_{j}}\|^{2}\right)$
19$\triangleright$ Inference Stage (Sec. 4.4)
20:Compute $\lambda_{i}\;,\;\lambda_{C_{1}}\;,\;\lambda_{G}$
21:$h_{i}^{\mathrm{final}}\leftarrow\lambda_{i}\,h_{i}\;+\;\lambda_{C_{1}}\,h_{C_{1}}\;+\;\lambda_{G}\,h_{G}$
22:$\hat{y}_{i}\leftarrow\sigma\Bigl{(}W_{F}\,h_{i}^{\mathrm{final}}\Bigr{)}$
@@ 4.5. Optimization and Training Details
    (20) $\mathcal{L}_{\text{sharp}}=\frac{1}{\lvert\mathcal{V}_{L}\rvert}\sum_{i\in\mathcal{V}_{\text{L}}}{1}\bigl{(}\hat{y}_{i}=y_{i}\bigr{)}\,u_{i}$
@@
The hyperparameters $\beta_{1}$ and $\beta_{2}$ trade off supervision against calibration, which can be tuned using a validation set. Alternatively, when a fully probabilistic treatment is desired, the last two terms may be replaced with a Bayesian regularizer such as an evidence lower bound (ELBO) or a PAC‑Bayesian bound.
@@
Initialization of Uncertainty.
For each node, we set the initial uncertainty $u_{i}^{(0)}$ according to the intrinsic reliability of its features. Concretely, we pre–compute a scalar confidence score with a node‑independent classifier $f_{\text{MLP}}(\cdot)$, a two-layer MLP with softmax trained on the raw attributes ${x}_{i}$. The initialization is given by:
(22) $u_{i}^{(0)}\;=\;1-\max f_{\text{MLP}}(x_{i})$
which means a node with highly discriminative features ($u^{(0)}_{i}\approx 0$) begin with low uncertainty score.
@@ 4.7. Computational Cost
Consider a graph $G$ with node feature dimension $d$, comprising $L$ local message‑passing layers and $M$ communities. The local layer first projects all $n$ node features, incurring a cost of $O(nd^{2})$, then computes attention and uncertainty weights over $m$ edges, each requiring $O(md)$. Subsequently, it aggregates neighbor messages at a complexity of $O(md)$ and updates node uncertainties, also in $O(md)$, leading to a per-layer computational cost of $O(nd^{2}+md)$. Combining the previously discussed costs, the local-layer computational complexity is given by $O(L(nd^{2}+md))$.
The community-level processing involves computing assignment scores for $n$ nodes across $M$ communities in $O(nMd)$, deriving $\arg\max$ assignments with complexity $O(nM)$, performing mean pooling and variance computation in $O(nMd)$, and finalizing aggregated community representations in $O(nd)$. Lastly, global integration and inference involve the global embedding with complexity $O(Md)$, computing global uncertainty in $O(Md)$, and combining final representations across all nodes at $O(nd)$, collectively amounting to $O(Md+nd)$. Thus, the overall computational complexity of HU-GNN is $O\left(L(nd^{2}+md)+nMd+nd\right)$.
@@ 5. Theoretical Analysis
We now present theoretical results that characterize the performance of HU-GNN. We focus on three aspects: generalization ability, convergence of the uncertainty propagation mechanism, and robustness in heterophilic graphs. Proof sketches or intuitions are provided for each subsection below.
@@ 5.1. PAC-Bayes Generalization Bounds
One of the advantages of modeling uncertainty is that it can improve generalization by avoiding overfitting to noisy signals. We derive a PAC-Bayesian generalization bound for HU-GNN, showing that it achieves a tighter dependency on graph complexity measures (such as node degrees) compared to a standard GNN. Our analysis is inspired by the PAC-Bayes bounds for GNNs (Liao et al., 2020), which showed that for a GCN (Kipf and Welling, 2016), the maximum node degree $\Delta_{\max}$ and weight norms control the generalization gap. Intuitively, high-degree nodes are problematic because they can aggregate many noisy signals, increasing variance.
@@
Theorem 1 (PAC-Bayes Generalization Bound for HU-GNN) Consider the HU-GNN model with $L$ layers (including hierarchical ones) trained on a graph $G$ for node classification. Let $\mathcal{D}$ be the full data distribution and $\mathcal{D}_{L}$ the training set (labels observed). Assume that $H[\cdot]$ stands for the (empirical) entropy of predictions. Given that the loss function is bounded, the following inequality holds with high probability for all posterior distributions $Q$ over the HU-GNN’s weights and a suitably chosen prior $P$:
@@
(23) $\begin{gathered}\mathbb{E}_{Q}[\text{Err}(\mathcal{D})]\;\leq\;\mathbb{E}_{Q}[\text{Err}(\mathcal{D}_{L})]\;+\;\\ \mathcal{O}\Big{(}\frac{1}{|\mathcal{D}_{L}|}\Big{[}\sum_{i\in\mathcal{D}_{L}}\min\{\tilde{\Delta}_{i},\Delta_{\max}\}\,+\,H[u_{L}]\Big{]}\Big{)}\;+\;\delta,\end{gathered}$
@@
where $\text{Err}(\mathcal{D}_{L})$ is training error. Notation $\tilde{\Delta}_{i}$ is an effective degree of node $i$ after uncertainty-based reweighting, $H[u_{L}]$ is a measure of uncertainty entropy across the graph, and symbol $\delta=\tilde{\mathcal{O}}\big{(}\frac{KL(Q\|P)+\log(1/\delta)}{|\mathcal{D}_{L}|}\big{)}$ is complexity term of order for confidence $1-\delta$. In particular, $\min\{\tilde{\Delta}_{i},\Delta_{\max}\}$ indicates that HU-GNN effectively caps the influence of high-degree nodes by reducing weights of edges from uncertain neighbors.
@@
The precise form of the bound is technical, but the key insight is that HU-GNN’s uncertainty mechanism leads to an effective degree $\tilde{\Delta}_{i}$, which is often much smaller than the raw degree $\Delta_{i}$. For example, if node $i$ has 100 neighbors but 90 of them are deemed highly uncertain, the effective degree in the model’s hypothesis space is closer to 10. This means the model is less likely to overfit based on those 90 noisy neighbors. As a result, our PAC-Bayesian bound does not blow up with the actual maximum degree of the graph, but rather with a lower quantity reflecting the filtered graph connectivity. This yields a tighter generalization bound compared to a standard GNN on the original graph (Liao et al., 2020). In addition, the newly added term $H[u_{L}]$ penalizes high uncertainty entropy. This makes the bound looser if the model remains very unsure (high entropy in uncertainties) across the graph. Minimizing this term essentially encourages the model to explain away uncertainty when possible, which aligns with training objectives that reduce uncertainty as confidence improves. The detailed proof uses a PAC-Bayes bound on a Gibbs classifier (Germain et al., 2009) that samples a random instantiation of the GNN weights, and leverages the convexity of the loss and the uncertainty gating to bound the change in loss when an edge weight is reduced. We also extend the perturbation analysis of message-passing networks to account for the random masking of edges by uncertainty, showing this acts like an $\ell_{0},\ell_{1}$ regularization on the adjacency. In conclusion, we prove that HU-GNN is theoretically justified to generalize better, especially on multiple benchmark graphs with noisy connections.